a proton moving horizontally at 3.5x10^4 m/s enters a uniform magnetic field of 0.075 t directed upward at 30°with the horizontal. find the magnitude of the magnetic force acting on the proton


1
Expert's answer
2022-04-15T16:40:11-0400

"F=qvB\sin\alpha=0.13\cdot 10^{-15}~N."